What is the slope of the line passing through points A(−18,24) and B(16,−8)?

Answer by math_helper(2461)


Slope = rise/run
Rise is the change in y values in traveling from one point to the other
(Run is the change in x values)
Often slope is represented by the letter m, for example in the slope-intercept form of a line y=mx+b
m = (-8-24)/(16-(-18)) = _____________________